\begin{thebibliography}{10}

\bibitem{BAIER2010879}
S.~Baier and M.~P. Young.
\newblock Mean values with cubic characters.
\newblock {\em Journal of Number Theory}, 130(4):879--903, 2010.

\bibitem{2017On}
L.~Bary-Soroker and P.~Meisner.
\newblock On the distribution of the rational points on cyclic covers in the absence of roots of unity.
\newblock {\em Mathematika}, 2017.

\bibitem{Boc}
S.~Bochner.
\newblock A theorem on analytic continuation of functions in several variables.
\newblock {\em Ann. of Math. (2)}, 39(1):14--19.

\bibitem{Cech1}
M.~{\v C}ech.
\newblock The {R}atios conjecture for real {D}irichlet characters and multiple {D}irichlet series.
\newblock {\em Trans. Amer. Math. Soc.}, 377:3487--3528.

\bibitem{david2024non}
C.~David, A.~Faveri, A.~Dunn, and J.~Stucky.
\newblock Non-vanishing for cubic {H}ecke ${L}$-functions.
\newblock {\em arXiv preprint arXiv:2410.03048}, 2024.

\bibitem{2021Nonvanishing}
C.~David, A.~Florea, and M.~Lalin.
\newblock Nonvanishing for cubic ${L}$ -functions.
\newblock {\em Forum of Mathematics, Sigma}, 9, 2021.

\bibitem{david2025nonvanishing}
C.~David, A.~Florea, and M.~Lalin.
\newblock Nonvanishing of ${L}$--functions associated to fixed order characters over function fields.
\newblock {\em arXiv preprint arXiv:2506.07815}, 2025.

\bibitem{2019The}
C.~David, A.~Florea, and M.~Lalín.
\newblock The mean values of cubic ${L}$-functions over function fields.
\newblock {\em Algebra \&amp; Number Theory}, 2019.

\bibitem{GAO2024125}
P.~Gao and L.~Zhao.
\newblock First moment of central values of some primitive dirichlet ${L}$-functions with fixed order characters.
\newblock {\em Journal of Number Theory}, 261:125--142, 2024.

\bibitem{goel2025second}
S.~Goel and A.~Ray.
\newblock The second moment of cubic {D}irichlet ${L}$-functions over function fields.
\newblock {\em arXiv preprint arXiv:2505.12015}, 2025.

\bibitem{gulouglu2024mollified}
A.~G{\"u}lo{\u{g}}lu and H.~Yesilyurt.
\newblock Mollified moments of cubic {D}irichlet ${L}$-functions over the $e$isenstein field.
\newblock {\em Journal of Mathematical Analysis and Applications}, 533(2):128014, 2024.

\bibitem{hong2025mean}
Z.~Hong, Z.~Fang, and Z.~Zheng.
\newblock Mean value of cubic $ {L} $-funcitons with fixed genus.
\newblock {\em arXiv preprint arXiv:2503.17228}, 2025.

\bibitem{hong2025first}
Z.~Hong and Z.~Zheng.
\newblock The first moment of central value of primitive quartic $ {L} $-functions with fixed genus.
\newblock {\em arXiv preprint arXiv:2504.14291}, 2025.

\bibitem{Luo_2004}
W.~Luo.
\newblock On {H}ecke ${L}$-series associated with cubic characters.
\newblock {\em Compositio Mathematica}, 140(5):1191–1196, 2004.

\end{thebibliography}

\documentclass[11pt, reqno]{amsart}[11pt]
\usepackage{geometry}                % See geometry.pdf to learn the layout options. There are lots.
\usepackage{graphicx}
\usepackage{epstopdf}
\usepackage{xcolor}
\usepackage{soul}
\newtheorem{theorem}{Theorem}[section]
\newtheorem{remark}{Remark}[section]
\usepackage[commandnameprefix=always]{changes}
\usepackage{amsmath,amssymb,amscd,mathrsfs,tikz-cd,mathtools,scalerel,latexsym}
\usepackage{bm}%%%%处理参考文献
\usepackage{color}
\numberwithin{equation}{section}
\DeclareMathOperator{\H2}{\mathcal{H}_{q^2, g/2+1}}
\DeclareMathOperator{\A2}{\mathcal{A}_{q^2}}
\DeclareMathOperator{\Hq}{\mathcal{H}_{q^2}}
\DeclareMathOperator{\Aq}{\mathcal{A}_{q}}
\DeclareMathOperator*{\Res}{Res}  
\newtheorem{lemma}{Lemma}[section]
\DeclareMathOperator{\moda}{mod}
\newtheorem{definition}{Definition}[section]

\title{Twisted second moment of primitive cubic $L$-functions}

\author{Ziwei Hong}
\address{School of Mathematical Sciences, Renmin University of China, Beijing, P.R. China}
\email{hongziwei@live.com}

\author{Zhiyong Zheng}
\address{School of Mathematics, 
Renmin University of China, Beijing, P.R. China}
\email{zhengzy@ruc.edu.cn}

\date{July 2025}

\begin{document}

\begin{abstract}
We investigate the mean value of the twisted second moment of primitive cubic $L$-functions over $\mathbb{F}_q(T)$ in the non-Kummer setting. Specifically, we study the sum
 \begin{equation*}
 \sum_{\substack{\chi\ primitive\ cubic\\ genus(\chi)=g}}\chi(h_1)\bar{\chi}(h_2)|L_q(\frac{1}{2}, \chi)|^2,
 \end{equation*}
where $L_q(s,\chi)$ denotes the $L$-function associated with primitive cubic character $\chi$. Employing a double Dirichlet series approach, we establish an error term of size $q^{(\frac{1}{2}+\varepsilon)g}$. 
\end{abstract}

\maketitle
\noindent {\bf Mathematics Subject Classification (2020)}: 11M06, 11M41, 11N37, 11L05, 11L40   

\noindent {\bf Keywords}:  central values, cubic $L$-functions, twisted moment, function field

\section{Introduction}

In this paper, we study the twisted second moment of primitive cubic Dirichlet $L$-functions $L_q(s,\chi)$ evaluated at the central point $s=\frac{1}{2}$, where $\chi$ ranges over primitive cubic characters of $\mathbb{F}_q[T]$ of fixed genus $g$, under the assumption that $q \equiv 2 \pmod{3}$.

 \subsection{Background}
The study of moments of $L$-functions is a central topic in analytic number theory, offering insight into the value distribution, non-vanishing, and zero statistics of these functions across various families. 

In the number field setting, Luo \cite{Luo_2004} investigated the first moment of cubic Dirichlet twists over $\mathbb{Q}(\xi_3)$, while Baier and Young \cite{BAIER2010879} considered the first moment of central values of Dirichlet $L$-functions associated with primitive cubic characters over $\mathbb{Q}$. G\"uloglu and Yesilyurt \cite{gulouglu2024mollified} studied the mollified first moment over the Eisenstein field, and also provided upper bounds for the mollified second moment. More recently, David et al. \cite{david2024non} analyzed the mollified second moment in the same setting and obtained a power-saving error term, which they used to study the non-vanishing of cubic $L$-functions.

In the function field context, significant progress has been made. David, Florea and Lalín \cite{2019The} computed the first moment of cubic L-functions over function fields with characters of fixed genus, assuming $q \equiv 2 \pmod{3}$. Later, in \cite{2021Nonvanishing}, they established upper bounds for a mollified second moment to study non-vanishing at the central point. Specifically, they proved that
\begin{align}\label{e}
    \sum_{\chi\in\mathcal{C}(g)}|L(\frac{1}{2},\chi)|^k|M(\chi;\frac{1}{\kappa})|^{k\kappa}\ll q^g.
\end{align}
where $\mathcal{C}(g)$ denotes the set of primitive cubic characters of genus $g$ and $M(\chi;\frac{1}{\kappa})$ is a Dirichlet polynomial. In their most recent work \cite{david2025nonvanishing}, they applied the theory of low-lying zeros to further investigate non-vanishing phenomena. In a complementary direction, Goel and Ray \cite{goel2025second} computed the (untwisted) second moment of cubic $L$-functions and showed that
\begin{align}
    \sum_{\substack{\chi~primitive~cubic\\ genus(\chi)=g}}L_{q}(\frac{1}{2},\chi)^2=\frac{g(g+2)A_q(\frac{1}{q^2},\frac{1}{q^{3/2}})\zeta_q(3/2)^2}{8\zeta_q(3)}q^{g+2}+O(q^g),
\end{align}
where $A_q$ is an explicitly defined function. Their method builds upon the framework developed in \cite{2019The}.

Despite these advances, the asymptotic of twisted moments in the cubic case has not yet been studied.

\subsection{Our contribution} In this paper, we address this gap and establish the first result in this direction. Specifically, we consider the twisted second moment in the non-Kummer setting. Our main theorem is as follows:
\begin{theorem}
Let $q$ be an odd prime power such that $q\equiv 2\moda 3$ and $h_1, h_2\in\mathbb{F}_q[T]$ be fixed polynomials. Then, for arbitrary $\varepsilon>0$, 
\begin{align}\label{main}
   \sum_{\substack{\chi\ primitive\ cubic\\ genus(\chi)=g}}\chi(h_1)\bar{\chi}(h_2)|L_q(\frac{1}{2}, \chi)|^2=(1-q^{2})SC(h_1,h_2)P(q^{-2})q^g+O(q^{(\frac{1}{2}+\varepsilon)g}),    
\end{align}
 where $P(u)$, $S$ and $C(h_1,h_2)$ are given in Section~\ref{residue}.
\end{theorem}
Here $P(u)$ is a function defined by \eqref{P}, $S$ is a constant from \eqref{S} and $C(h_1,h_2)$ is a constant depending on the twisting polynomials $h_1$ and $h_2$. $L_q(s,\chi)$ is the $L$-function associeted to $\chi$.

Our result is consistent with the bound in \eqref{e}, and provides new quantitative information about the second moment with nontrivial twists.

\subsection{Methodology}

A key innovation in our approach is the use of double Dirichlet series. Using Perron’s formula, we express the twisted second moment as a contour integral involving such a series. By representing the double Dirichlet series in different ways, we identify multiple regions of convergence. Basic tools from complex analysis imply that the convex hull of these regions also lies within the domain of absolute convergence, allowing us to shift contours further and isolate the main term.

This method—originating from studies of moments of L-functions of higher-order characters (see \cite{GAO2024125, hong2025first,hong2025mean})—proves to be powerful and flexible. In particular, it avoids the intricate analysis of the distribution of cubic Gauss sums, which significantly simplifies the computation and reduces technical complexity.

Compared to the approaches used in \cite{2019The} and \cite{goel2025second}, our method is more direct and computationally efficient.

\section{Preliminary}

In this section, we recall some necessary well-known results on primitive cubic characters and $L$-functions. In fact, all these results can be found in \cite{2019The}.

We work in the non-Kummer setting, where q is an odd prime power such that $q \equiv 2 \moda{3}$. For simplicity, we do not consider the corresponding results in the Kummer setting ($q \equiv 1 \moda{3}$), although a similar asymptotic formula for the first moment of cubic L-functions in that case can be derived through analogous methods.
%The foundational details can be found in Section 2 of \cite{2019The}, and we summarize the necessary results here.

%We advise readers to skip this section initially and proceed directly to Section~3, consulting the lemmas herein only when they are explicitly referenced in the computational steps of Section~3.

\subsection{Primitive cubic characters and $L$-functions}
In the non-Kummer setting, defining primitive cubic characters over $\mathbb{F}_q[T]$ is more intricate. Following the approach of \cite{2017On}, it is natural to restrict a cubic character defined over $\mathbb{F}_{q^2}[T]$ to obtain a cubic character over $\mathbb{F}_q[T]$.

We fix once and for all an isomorphism $\Omega$ between the cubic roots of $1$ in $\mathbb{C}^*$ and the cubic roots of $1$ in $\mathbb{F}^*_{q^2}$. We define the cubic residue symbol $\chi_{\pi}$, for $\pi$ an irreducible monic polynomial in $\mathbb{F}_{q^2}[T]$ such that $\pi\pi^{\sigma}=P$ is an irreducible polynomial in $\mathbb{F}_q[T]$, where $\sigma$ is the generator of $Gal(\mathbb{F}_{q^2}/\mathbb{F}_q)$. Let $a\in\mathbb{F}_{q^2}[T]$. If $\pi |a$, then $\chi_{\pi}(a)=0$, and otherwise $\chi_{\pi}(a)=\alpha$, where $\alpha$ is the unique root of unity in $\mathbb{C}$ such that 
\begin{equation*}
    a^{\frac{q^{2\deg\pi}-1}{3}}\equiv\Omega(\alpha)\moda \pi.
\end{equation*}
Then $\chi_{\pi}|_{\mathbb{F}_q[T]}$ is a cubic character module $P$.

We extend the definition by multiplicity. For any monic polynomial $F\in\mathbb{F}_{q^2}[T]$, $F=\pi_1^{e_1}\dots\pi_s^{e_s}$ with distinct $\pi_i$, we define $\chi_F=\chi_{\pi_1}^{e_1}\dots\chi_{\pi_s}^{e_s}$ . We restrict $\chi_{F}$ to $\mathbb{F}_q[T]$, then $\chi_{F}$ is a cubic character on $\mathbb{F}_q[T]$. $\chi_F$ is primitive if and only if $e_i=1$ for i=1,2,...,s and $F$ has no divisor in $\mathbb{F}_q[T]$. It it is so, then $\chi_F$ has conductor $FF^{\sigma}$.

 Let $\mathcal{A}_q$ denote the set of monic polynomials over $\mathbb{F}_q$. Let $\H2$ represent the set of monic, square-free polynomials in $\A2$ of degree $g/2+1$. With above notations, in \cite{2019The}, David et al show that
 \begin{lemma}
      Suppose $q\equiv 2\moda 3$, Then,
     \begin{equation}\label{equation}
 \sum_{\substack{\chi\ primitive\ cubic\\ genus(\chi)=g}}\chi(h_1)\bar{\chi}(h_2)|L_q(\frac{1}{2}, \chi)|^2=  \sum_{\substack{F\in\H2 \\ P|F\Rightarrow P\not\in \mathbb{F}_q[t]}}\chi_F(h_1)\bar{\chi}_F(h_2)|L_q(\frac{1}{2}, \chi_{F})|^2.
 \end{equation}
 \end{lemma}
 As 
 \begin{equation*}
     \chi_F(\alpha)=\Omega^{-1}\left(\alpha^{\frac{q^{2\deg F}-1}{3}}\right)
 \end{equation*}
 for $\alpha\in\mathbb{F}_q\subset\mathbb{F}_{q^2}$, and q is odd and $q\equiv 2\moda 3$, we remark that all cubic characters over $\mathbb{F}_q[T]$ are even. Hence, we have the following functional equation for even $L$-functions:

 \begin{lemma}[Functional equation] \label{fe}
     Let $F\in\mathcal{H}_{q^2}$ and suppose $F$ has no divisor in $\Aq$. Then
     $$L_q(s,\chi_{F})=\epsilon(\chi_F)q^{2s-1}\frac{1-q^{-s}}{1-q^{s-1}}\frac{L_q(1-s,\overline{\chi_F})}{|F|_2^{s-\frac{1}{2}}},$$
     where $\epsilon(\chi_F)=q^{-\deg F}G(\chi_F)$, $G(\chi_F)$ is the Gauss sum.
 \end{lemma}
 Here $|F|_2=q^{2\deg F}$ stands for the modulus of polynomials in $\mathbb{F}_{q^2}[T]$. And the Gauss sum is defined as
  \begin{definition}
     For $\chi$ a primitive character of the modulus h on $\mathbb{F}_q[T]$, let
     $$G_q(\chi)=\sum_{a \moda h}\chi(a)e_q(\frac{a}{h}).$$
 \end{definition}

Next, we present the upper bound of $L$-functions. The folllowing lemma restates Lemma 2.6 and Lemma 2.7 from \cite{2019The}.
 \begin{lemma}[Lindel\"of Hypothesis]
 \label{LLH}
     Let $\chi$ be a primitive cubic character of conductor $h$ defined over $\mathbb{F}_q[T]$. Then, for $\Re(s)\ge\frac{1}{2}$ and all $\varepsilon>0$,
     \begin{align}
         |L_q(s,\chi)|\ll q^{\varepsilon\deg h};
     \end{align}
     for $\Re(s)\ge 1$ and for all $\varepsilon>0$
     \begin{align}
         |L_q(s,\chi)|\gg q^{-\varepsilon\deg h}.
     \end{align}
 \end{lemma}

 \subsection{Multivarible Complex Analysis Lemmas}
Our approach relies on two foundational results from multivariable complex analysis. We begin by introducing the concept of a tube domain.
\begin{definition}
		An open set $T\subset\mathbb{C}^n$ is a tube if there is an open set $U\subset\mathbb{R}^n$ such that $T=\{z\in\mathbb{C}^n:\ \Re(z)\in U\}.$
\end{definition}
	
   For any set $U\subset\mathbb{C}^n$, we define $T(U)=U+i\mathbb{R}^n\subset \mathbb{C}^n$.  We quote the following Bochner's Tube Theorem \cite{Boc}.
\begin{theorem}
\label{Bochner}
		Let $U\subset\mathbb{R}^n$ be a connected open set and $f(z)$ be a function holomorphic on $T(U)$. Then $f(z)$ has a holomorphic continuation to
the convex hull of $T(U)$.
\end{theorem}

 We denote the convex hull of an open set $T\subset\mathbb{C}^n$ by $\widehat T$.  Our next result is \cite[Proposition C.5]{Cech1} on the modulus of holomorphic continuations of multivariable complex functions.
\begin{theorem}

\label{Extending inequalities}
		Assume that $T\subset \mathbb{C}^n$ is a tube domain, $g,h:T\rightarrow \mathbb{C}$ are holomorphic functions, and let $\tilde g,\tilde h$ be their
holomorphic continuations to $\widehat T$. If  $|g(z)|\leq |h(z)|$ for all $z\in T$ and $h(z)$ is nonzero in $T$, then also $|\tilde g(z)|\leq
|\tilde h(z)|$ for all $z\in \widehat T$.
\end{theorem}

\section{Analytical behavior of $A_2(u,v)$}
In this section, we analyze the twisted second moment by studying the analytic behavior of the double Dirichlet series $A_2(u,v)$.
 Using Perron's formula for function fields, we express a sum of coefficients as:
 \begin{equation}
     \sum_{n\le N}a(n)=\frac{1}{2\pi i}\oint_{|u|=r}(\sum_{n=0}^{\infty}a(n)u^n)\frac{du}{(1-u)u^{N+1}}.
 \end{equation}
This allows us to rewrite \eqref{equation} as 
\begin{align}
    \sum_{\substack{F\in\H2 \\ P|F\Rightarrow P\not\in \mathbb{F}_q[t]}}\chi_F(h_1)\bar{\chi}_F(h_2)|L_q(\frac{1}{2}, \chi_{F})|^2=\frac{1}{2\pi i}\oint_{|u|=r}A_2(u,q^{-\frac{1}{2}})\frac{du}{u^{g/2+2}},
\end{align}
where $A_2(u,v)$ is derived from the double Dirichlet series
\begin{align}\label{A2}
    A_2(s,w)=\sum_{\substack{F\in\Hq\\ P|F\to P\not\in\Aq}}\frac{\chi_F(h_1)\overline{\chi_F(h_2)}|L_q(w,\chi_F)|^2}{|F|_2^s}
\end{align}
by substituting $u=q^{-2s}$ and $v=q^{-w}$.

Hereafter, we will use $A_3(s,w)$ and $A_3(u,w)$, $L_q(w,\chi)$ and $L_q(v,\chi)$ interchangeably when there is no ambiguity.

To understand the behavior of $A_2(u,v)$, we reformulate this expression to identify its domain of convergence and any singularities. We also derive an appropriate bounding function.

\subsection{First convergence region}\label{main tech}
To remove the condition that F has no prime divisor in $\Aq$, we apply M\"obius inversion 
 \begin{equation}     
 \sum_{\substack{D\in\Aq\\ D|F}}\mu(D)=\left\{ \begin{array}{ll}
         1 & \mbox{$F$ has no prime divisor in } \mathbb{F}_q[T]; \\
         0 & \mbox{otherwise}.
     \end{array}
     \right.
 \end{equation}
Substituting this into our expression for $A_2(s,w)$, we get
\begin{align*}
    A_2(s,w)=&\sum_{\substack{F\in\Hq\\ P|F\to P\not\in\Aq}}\frac{\chi_F(h_1)\overline{\chi_F(h_2)}|L(w,\chi_F)|^2}{|F|_2^s}\\
    =&\sum_{D\in\Aq}\mu(D)\sum_{\substack{F\in\Hq\\ (F,D)=1}}\frac{\chi_{DF}(h_1)\overline{\chi_{DF}(h_2)}|L(w,\chi_{DF})|^2}{|DF|_2^s}\\
    =&\sum_{D\in\Aq}\frac{\mu(D)}{|D|_2^s}\sum_{\substack{F\in\Hq\\ (F,D)=1}}\frac{\chi_{DF}(h_1)\overline{\chi_{DF}(h_2)}}{|F|_2^s}\sum_{N_1,N_2\in\Aq}\frac{\chi_{DF}(N_1)\overline{\chi_{DF}(N_2)}}{|N_1|^w|N_2|^{\bar{w}}}\\
    =&\sum_{N_1,N_2\in\Aq}\frac{1}{|N_1|^w|N_2|^{\bar{w}}}\sum_{D\in\Aq}\frac{\mu(D)\chi_D(h_1N_1)\overline{\chi_D(h_2N_2)}}{|D|_2^s}\sum_{\substack{F\in\Hq\\ (F,D)=1}}\frac{\chi_{F}(h_1N_1)\overline{\chi_{F}(h_2N_2)}}{|F|_2^s}.
\end{align*}
Here $|N|=q^{-\deg N}$ denotes the modulus of polynomial in $\Aq$. From now on, we use $P_1$ to denote prime in $\Aq$ and use $P_2$ to denote prime in $\A2$. Since cubic characters satisfy $\chi^2 = \overline{\chi}$, we simplify
\begin{align*}
    =&\sum_{N_1,N_2\in\Aq}\frac{1}{|N_1|^w|N_2|^{\bar{w}}}\sum_{D\in\Aq}\frac{\mu(D)\chi_D(h_1h_2^2N_1N_2^2)}{|D|_2^s}\sum_{\substack{F\in\Hq\\ (F,D)=1}}\frac{\chi_{F}(h_1h_2^2N_1N_2^2)}{|F|_2^s}\\
    =&\sum_{N_1,N_2\in\Aq}\frac{1}{|N_1|^w|N_2|^{\bar{w}}}\sum_{D\in\Aq}\frac{\mu(D)\chi_D(h_1h_2^2N_1N_2^2)}{|D|_2^s}\prod_{\substack{P_2\in\A2\\ P_2\nmid D}}\left(1+\frac{\chi_{P_2}(h_1h_2^2N_1N_2^2)}{|P_2|_2^s}\right)\\
    =&\sum_{N_1,N_2\in\Aq}\frac{1}{|N_1|^w|N_2|^{\bar{w}}}\frac{L_{q^2}(s,\chi^{(hN)})}{L_{q^2}(2s,\overline{\chi^{(hN)}})}\sum_{D\in\Aq}\frac{\mu(D)\chi_D(hN)}{|D|_2^s}\prod_{\substack{P_2\in\A2\\ P_2| D}}\left(1+\frac{\chi_{P_2}(hN)}{|P_2|_2^s}\right)^{-1}.
\end{align*}
Here we take $h_1h_2^2N_1N_2^2=hN$ for simplicity. $\chi^{(hN)}(F)=\chi_F(hN)$ is a Hecke character.
 The second equality in above equation uses the trick
 \begin{align*}
     &\prod_{P_2\nmid D}\left(1+\frac{\chi_{P_2}(h_1h_2^2N_1N_2^2)}{|P_2|_2^s}\right)\\
     =&\prod_{P_2\nmid D}\frac{1-\frac{\overline{\chi}_{P_2}(h_1h_2^2N_1N_2^2)}{|P_2|_2^{2s}}}{1-\frac{\chi_{P_2}(h_1h_2^2N_1N_2^2)}{|P_2|_2^s}}=\frac{\prod\limits_{P_2}\left(1-\frac{\chi_{P_2}(h_1h_2^2N_1N_2^2)}{|P_2|_2^s}\right)^{-1}}{\prod\limits_{P_2}\left(1-\frac{\overline{\chi}_{P_2}(h_1h_2^2N_1N_2^2)}{|P_2|_2^{2s}}\right)^{-1}}\prod_{P_2|D}\left(1+\frac{\chi_{P_2}(h_1h_2^2N_1N_2^2)}{|P_2|_2^s}\right)^{-1}\\
     =&\frac{L_{q^2}(s,\chi^{(hN)})}{L_{q^2}(2s,\overline{\chi^{(hN)}})}\prod_{P_2|D}\left(1+\frac{\chi_{P_2}(h_1h_2^2N_1N_2^2)}{|P_2|_2^s}\right)^{-1}.
 \end{align*}
   We now use the following lemma to simplify the inner sumof $A_2(s,w)$.
 \begin{lemma}
     If $D, F\in\Aq$ and $(D, F)=1$, then we have $\chi_D(F)=1$.
 \end{lemma}
 \begin{proof}
     Let $\sigma: \mathbb{F}_{q^2}\to \mathbb{F}_{q^2}$ be the non-trivial Frobenius automorphism. 
     We suppose $D=P$ is irreducible. By definition, we have
     $$F^{\frac{q^{2\deg P}-1}{3}}\equiv \Omega(\alpha) \moda P,$$
     which is equivalent to say
     $F^{\frac{q^{2\deg P}-1}{3}}=f(T)P+\Omega(\alpha).$ Then $\sigma$ gives  $F^{\frac{q^{2\deg P}-1}{3}}=f^{\sigma}(T)P+\sigma(\Omega(\alpha))\equiv \Omega(\alpha) \moda P$. Note the fact that if two constants are equivalent module $P$, then the two constants are the same. We finally have $\sigma(\Omega(\alpha))= \Omega(\alpha)$. Hence $\Omega(\alpha)=1$. 
     For $D=P_1^{e_1}\cdots P_k^{e_k}$, we have $\chi_D=\chi_{P_1}^{e_1}\cdots \chi_{P_k}^{e_k}$. This complete the proof.
 \end{proof}
 Applying the lemma, we reduce
 \begin{align*}
     A_2(s,w) =&\sum_{N_1,N_2\in\Aq}\frac{1}{|N_1|^w|N_2|^{\bar{w}}}\frac{L_{q^2}(s,\chi^(hN))}{L_{q^2}(2s,\overline{\chi^{(hN)}})}\sum_{\substack{D\in\Aq\\ (D,hN)=1}}\frac{\mu(D)}{|D|_2^s}\prod_{\substack{P_2\in\A2\\ P_2| D}}\left(1+\frac{\chi_{P_2}(hN)}{|P_2|_2^s}\right)^{-1}\\
      =&\sum_{N_1,N_2\in\Aq}\frac{1}{|N_1|^w|N_2|^{\bar{w}}}\frac{L_{q^2}(s,\chi^(hN))}{L_{q^2}(2s,\overline{\chi^{(hN)}})}\prod_{\substack{P_1\in\Aq\\ P_1\nmid hN}}\left(1-\frac{1}{|P_1|_2^s}\prod_{P_2|P_1}\left(1+\frac{\chi_{P_2}(hN)}{|P_2|_2^s}\right)^{-1}\right)\\
      =&\sum_{N_1,N_2\in\Aq}\frac{1}{|N_1|^w|N_2|^{\bar{w}}}\frac{L_{q^2}(s,\chi^(hN))}{L_{q^2}(2s,\overline{\chi^{(hN)}})}P(s,\chi^{(hN)})\prod_{\substack{P_1\in\Aq\\ P_1|hN}}\left(1-\frac{1}{|P_1|_2^s}\right)^{-1},
 \end{align*}
 where
 \begin{align*}
     P(s,\chi^{(hN)})=\prod_{P_1\in\Aq}\left(1-\frac{1}{|P_1|_2^s}\prod_{P_2|P_1}\left(1+\frac{\chi_{P_2}(hN)}{|P_2|_2^s}\right)^{-1}\right).
 \end{align*}
For $\Re(s)>\frac{1}{2}$, we have
\begin{align}\label{P1}
    P(s,\chi^{(N)})=\zeta_q^{-1}(2s)\prod\limits_{P_1\in\Aq}\left(1+O(\frac{1}{|P_2|^{3s}_2})\right),
\end{align}
where $\zeta_q(s)=\frac{1}{1-q^{1-s}}$. As $(1-\frac{1}{|P_1|_2^s})^{-1}=1+\frac{1}{|P_1|_2^s-1}\ll |P_1|_2^{\max\{0,-\Re(s)\}}$, we will obtain 
\begin{align}\label{P2}
 \prod_{\substack{P_1|hN}}\left(1-\frac{1}{|P_1|_2^s}\right)^{-1}\ll |hN|_2^{\max\{0,-\Re(s)\}+\varepsilon}.  
\end{align}
 
 Substituting $u=q^{-2s}$, $v=q^{-w}$, we finally express
 \begin{align}\label{A2S1}
     A_2(u,v)=\sum_{N_1,N_2\in\Aq}v^{\deg N_1}\bar{v}^{\deg N_2}\frac{L_{q^2}(u,\chi^{(hN)})}{L_{q^2}(u^2,\overline{\chi^{(hN)}})}P(u,\chi^{(hN)})\prod_{\substack{P_1\in\Aq\\ P_1|hN}}\left(1-u^{\deg P_1}\right)^{-1},
 \end{align}
where $L_{q^2}(u,\chi)$ is the $L$-function associated $\chi$ over $\mathbb{F}_{q^2}(T)$ and 
\begin{align*}
    P(u,\chi^{(hN)})=\prod_{P_1}\left(1-u^{\deg P_1}\prod_{P_2|P_1}\left(1+\chi_{P_2}(hN)u^{\deg P_2}\right)^{-1}\right).
\end{align*}
 The ratio $\frac{L_{q^2}(u,\chi^{(hN)})}{L_{q^2}(u^2,\overline{\chi^{(hN)}})}$ is a rational function in $u$. For $|u|<q^{-1}$ and arbitrary $\varepsilon>0$, Lemma~\ref{LLH}(Lindel\"of Hypothesis) gives 
\begin{align}\label{eL}
    \frac{L_{q^2}(u,\chi^{(hN)})}{L_{q^2}(u^2,\overline{\chi^{(hN)}})}\ll |hN|_2^{\varepsilon}.
\end{align}
Combing \eqref{P1}, \eqref{P2}, \eqref{A2S1} and \eqref{eL}, we see that $(u-q^{-2})A_2(u,v)$ has the bound
\begin{align}
    (u-q^{-2})A_2(u,v)\ll |u-q^{-2}||h_1h_2^2|_2^{\varepsilon}\sum_{N_1}|q^{2\varepsilon}v|^{\deg N_1}\sum_{N_2}|q^{4\varepsilon}v|^{\deg N_2}
\end{align}
and the right-hand side series is convergent in 
\begin{align}
    S_1=\{(u,v)||u|<q^{-1}, |v|<q^{-1}\},
\end{align}
except the possible simple pole at $u=q^{-2}$.

  \subsection{Residue at $u=q^{-2}$}\label{residue}
  From \eqref{A2S1}, $A_2(u,v)$ has a possible simple pole at $u=q^{-2}$. This pole arises when the product $h_1h_2^2N_1N_2^2$ forms a perfect cube. We now compute the residue of $A_2(u,v)$ at $u=q^{-2}$.\small{
  \begin{align*}
      &\Res_{u=q^{-2}}A(u,q^{-\frac{1}{2}})\\
      =& \Res_{u=q^{-2}}\sum_{\substack{N_1,N_2\in\Aq\\h_1h_2^2N_1N_2^2~ is~ a~ cube}}\frac{1}{|N_1N_2|^{\frac{1}{2}}}\frac{\zeta_{q^2}(u)}{\zeta_{q^2}(u^2)}\prod_{\substack{P_2\in\A2\\ P_2|hN}}\left(1+u^{\deg P_2}\right)^{-1}\prod_{\substack{P_1\in\Aq\\ P_1|hN}}\left(1-u^{\deg P_1}\prod_{P_2|P_1}(1+u^{\deg P_2})^{-1}\right)^{-1}P(u)\\
      =&\sum_{\substack{N_1,N_2\\h_1h_2^2N_1N_2^2~ is~ a~ cube}}\frac{1}{|N_1N_2|^{\frac{1}{2}}}\prod_{\substack{P_2\in\A2\\ P_2|hN}}\left(1+u^{\deg P_2}\right)^{-1}\prod_{\substack{P_1\in\Aq\\ P_1|hN}}\left(1-u^{\deg P_1}\prod_{P_2|P_1}(1+u^{\deg P_2})^{-1}\right)^{-1}P(q^{-2})\Res_{u=q^{-2}}\frac{\zeta_{q^2}(u)}{\zeta_{q^2}(u^2)}\\
      =&(q^{-4}-q^{-2})S(q^{-\frac{1}{2}})P(q^{-2}),
  \end{align*}}
where
\begin{align}
    S(v)&=\sum_{\substack{N_1,N_2\\h_1h_2^2N_1N_2^2~ is~ a~ cube}}v^{\deg N_1N_2}\prod_{\substack{P_2\in\A2\\ P_2|hN}}\left(1+\frac{1}{|P_2|_2}\right)^{-1}\prod_{\substack{P_1\in\Aq\\ P_1|hN}}\left(1-\frac{1}{|P_1|_2}\prod_{P_2|P_1}(1+\frac{1}{|P_2|_2})^{-1}\right)^{-1},
    %&=\sum_{\substack{N_1,N_2\\h_1h_2^2N_1N_2^2~ is~ a~ cube}}\frac{1}{|N_1N_2|^{\frac{1}{2}}}\prod_{P_1|hN}\left(1+\frac{2}{|P_1|}\right)^{-1}\notag\\
\end{align}
and
\begin{align}\label{P}
    P(u)=\prod_{\substack{P_1\in\Aq}}\left(1-u^{\deg P_1}\prod_{P_2|P_1}(1+u^{\deg P_2})^{-1}\right).
\end{align}
For each prime $P\in\Aq$, we define
\begin{align}
    ord_P(h_1)=a_p;\\
    ord_P(h_2)=b_p;\\
    ord_P(N_1)=c_p;\\
    ord_P(N_2)=d_p.
\end{align}
Next, we rewrite $S(v)$ as an Eular product

\begin{align*}
    S(v)=\prod_{\substack{P\in\Aq\\ P|h}}\left(\prod_{\substack{ P_2|P}}\left(1+\frac{1}{|P_2|_2}\right)^{-1}\left(1-\frac{1}{|P|_2}\prod_{P_2|P}(1+\frac{1}{|P_2|_2})^{-1}\right)^{-1}\sum_{3|a_P+2b_P+c_P+2d_P}v^{\deg P(c_P+d_P)}\right)\\
    \times\prod_{\substack{P\in\Aq\\ P\nmid h}}\left(1+\prod_{\substack{ P_2|P}}\left(1+\frac{1}{|P_2|_2}\right)^{-1}\left(1-\frac{1}{|P|_2}\prod_{P_2|P}(1+\frac{1}{|P_2|_2})^{-1}\right)^{-1}\sum_{3|c_P+2d_P}v^{\deg P(c_P+d_P)}\right) \\
%    =&\prod_{P}G_P(v) C(h_1,h_2) 
\end{align*}
and separate the factors related to $h_1$,$h_2$ 
\begin{align*}
    S(v)=\prod_{P}G_P(v) C(h_1,h_2),
\end{align*}
where
\begin{align}
    G_P(v)=1+\prod_{\substack{ P_2|P}}\left(1+\frac{1}{|P_2|_2}\right)^{-1}\left(1-\frac{1}{|P|_2}\prod_{P_2|P}(1+\frac{1}{|P_2|_2})^{-1}\right)^{-1}\sum_{3|c_P+2d_P}v^{\deg P(c_P+d_P)}
\end{align}
and
\begin{align}
    C(h_1,h_2,v)=\prod_{P|h}\frac{\prod\limits_{\substack{ P_2|P}}\left(1+\frac{1}{|P_2|_2}\right)^{-1}\left(1-\frac{1}{|P|_2}\prod\limits_{P_2|P}(1+\frac{1}{|P_2|_2})^{-1}\right)^{-1}\sum\limits_{3|a_P+2b_P+c_P+2d_P}v^{\deg P(c_P+d_P)}}{G_P(v)}.
\end{align}
For each prime $P\in\Aq$, we distinguish two cases depending on whether $\deg P$ is even or odd (noting that $P\in\Aq$ splits in $\A2$ if and only if $\deg P$ is even). In fact, we have
\begin{align*}
    G_P(v)=\begin{cases}
    1+\sum\limits_{3|c_P+2d_P}v^{\deg P(c_P+d_P)},& \mbox{if } 2\nmid\deg P;\\
        1+\left(1+\frac{2}{|P|}\right)^{-1}\sum\limits_{3|c_P+2d_P}v^{\deg P(c_P+d_P)}, &\mbox{otherwise.}
    \end{cases}
\end{align*}

Using the orthogonality of primitive cubic root
\begin{align}
    \frac{1}{3}\sum_{k=0}^2\varpi^{kn}=\begin{cases}
    1, &n\equiv 0\moda 3;\\
    0, &otherwise,
    \end{cases}
\end{align}
for $2|\deg P$, we have
\begin{align*}
    G_P(v)&=1+(1+\frac{2}{|P|})^{-1}\sum_{3|c_P+2d_P}v^{\deg P(c_P+d_P)}\\
    &=1+(1+\frac{2}{|P|})^{-1}(\sum_{\substack{c_P\ge 0,d_P\ge 0\\ }}v^{\deg P(c_P+d_P)}\frac{1}{3}\sum_{k=0}^2\varpi^{k(c_P+2d_P)}-1)\\
    &=1+(1+\frac{2}{|P|})^{-1}(\frac{1}{3}\sum_{k=0}^2\sum_{c_P,d_P}\varpi^{kc_P}v^{\deg Pc_P}\varpi^{2kd_P}v^{\deg Pd_P}-1)\\
    &=1+(1+\frac{2}{|P|})^{-1}(\frac{1}{3}\sum_{k=0}^2\frac{1}{1-\varpi^{k}v^{\deg P}}\frac{1}{1-\varpi^{2k}v^{\deg P}}-1)\\
    &=1+(1+\frac{2}{|P|})^{-1}\frac{v^{2\deg P}(1+v^{\deg P}-v^{2\deg P})}{(1-v^{2\deg P})^2(1+v^{\deg P}+v^{2\deg P})}.
\end{align*}
For $2\nmid\deg P$, we have
\begin{align*}
    G_P(v)=&1+\sum\limits_{3|c_P+2d_P}v^{\deg P(c_P+d_P)}\\
    =&\frac{1}{3}\sum_{k=0}^2(\sum_{c_P\ge 0}v^{\deg Pc_P}\varpi^{kc_P})(\sum_{d_P\ge 0}v^{\deg Pc_P}\varpi^{2kd_P})\\
    =&\frac{1}{3}\sum_{k=0}^2\frac{1}{1-\varpi^kv^{\deg P}}\frac{1}{1-\varpi^{2k}v^{\deg P}}\\
    =&\frac{1-v^{\deg P}+v^{2\deg P}}{(1-v^{\deg P})^2(1+v^{\deg P}+v^{2\deg P})}.
\end{align*}

Using similar method, we have
\begin{align*}    C(h_1,h_2,v)=\frac{\prod\limits_{P|h}\prod\limits_{\substack{ P_2|P}}\left(1+\frac{1}{|P_2|_2}\right)^{-1}\left(1-\frac{1}{|P|_2}\prod\limits_{P_2|P}(1+\frac{1}{|P_2|_2})^{-1}\right)^{-1}\frac{1}{3}(\frac{1}{(1-v^{\deg P})^2}+\frac{\varpi^{a_P+2b_P}+\varpi^{2a_P+b_P}}{1+v^{\deg P}+v^{2\deg P}})}{\prod\limits_{P|h}G_P(v)}.
\end{align*}

For simplicity, we denote 
\begin{align}\label{C}
    C(h_1,h_2)=C(h_1,h_2,q^{-\frac{1}{2}})
\end{align}
and
\begin{align}\label{S}
S=\prod_{P}G_P(q^{-\frac{1}{2}}).    
\end{align}
\begin{remark}
The factor $C(h_1, h_2)$ depends only on those primes dividing $h_1h_2$ for which $a_P + 2b_P \not\equiv 0 \mod 3$. The constant S is independent of $h_1$ and $h_2$.
\end{remark}
Hence, we have
\begin{align}
    \Res_{u=q^{-2}} A_2(u,q^{-\frac{1}{2}})=(q^{-4}-q^{-2})SC(h_1,h_2)P(q^{-2}).
\end{align}

\subsection{The second convergence region}
Lindel\"of Hypothesis gives
\begin{align*}
    |L_q(v,\chi_F)|\ll |F|_2^{\varepsilon}.
\end{align*}
Inserting this into $A_2(u,v)$ gives
\begin{align*}
    A_2(u,v)=&\sum_{\substack{F\in\Hq\\ P|F\to P\not\in\Aq}}\chi_F(h_1)\overline{\chi_F(h_2)}|L_q(v,\chi_F)|^2 u^{\deg F}\\
    &\ll \sum_{\substack{F\in\Hq\\ P|F\to P\not\in\Aq}} |F|_2^{\varepsilon}|u|^{\deg F} \\
    &\ll \sum_{\substack{F\in\Hq\\ P|F\to P\not\in\Aq}} |q^{2\varepsilon}u|^{\deg F}.
\end{align*}
Hence, $A_2(u,v)$ is convergent in region
\begin{align}
    S_{2,1}=\{(u,v)|~|v|\le q^{-\frac{1}{2}},|u|<q^{-2}\}.
\end{align}

By applying functional equation(Lemma~\ref{fe}) for $L_q(v,\chi_F)$ and again using the Lindel\"of Hypothesis(Lemma~\ref{LLH}), we can also write
\begin{align*}
    A_2(u,v)=&\sum_{\substack{F\in\Hq\\ P|F\to P\not\in\Aq}} \chi_F(h_1)\overline{\chi}_F(h_2)u^{\deg F}|\epsilon(\chi_F)|^2|(v^2q)^{-1}\frac{1-v}{1-\frac{1}{qv}}|^2|L_q(\frac{1}{qv},\chi_F)|^2|q^2v^4|^{\deg F}\\
    \ll & |(v^2q)^{-1}\frac{1-v}{1-\frac{1}{qv}}|^2\sum_{\substack{F\in\Hq\\ P|F\to P\not\in\Aq}}|q^{4\varepsilon+2}uv^4|^{\deg F}.
\end{align*}
The series on the right-hand side is convergent in
\begin{align}
    S_{2,2}=\{(u,v)||v|\ge q^{-\frac{1}{2}}, |uv^4|<q^{-3}\}.
\end{align}
%Then the second convergence region is
%\begin{align}
%    S_2=\mbox{convex hull of } S_{2,1}\cup S_{2,2}.
%\end{align}

\subsection{The third convergence region}
We now derive a third region of convergence by applying the functional equation, we express $A_2$ as
\begin{align*}
    A_2(u,v)=\big|\frac{1-v}{v^2q-v}\big|^2\sum_{\substack{F\in\Hq\\ P|F\to P\not\in\Aq}}\chi_F(h_1)\overline{\chi_F}(h_2)|L_q(\frac{1}{qv},\chi_F)|^2(q^2u|v^4|)^{\deg F}
\end{align*}
Applying the same analytic techniques used in Section~\ref{main tech}, we have
\begin{align}
    A_2(u,v)=\big|\frac{1-v}{v^2q-v}\big|^2\sum_{N_1,N_2\in\Aq}\frac{1}{(qv)^{\deg N_1}}\frac{1}{(q\bar{v})^{\deg N_2}}\frac{L_{q^2}(u|v|^4q^2,\chi^{(hN')})}{L_{q^2}((u|v|^4q^2)^2,\overline{\chi^{(hN')}})}\\
    \times P(u|v|^4q^2,\chi^{(hN')})\prod_{\substack{P_1\in\Aq\\ P_1|hN'}}\left(1-(q^2u|v^2|)^{\deg P_1}\right)^{-1},
\end{align}
here we take $hN'=h_1h_2^2N_1^2N_2$ for simplicity.

Lindel\"of Hypothesis shows, for $|uv^4q^2|>q^{-1}$,
\begin{align}
    |(u|v^4|q^{2}-q^{-2})A_2(u,v)|\ll |(u|v^4|q^{2}-q^{-2})|\big|\frac{1-v}{v^2q-v}\big|^2|h_1h_2^2|_2^{\varepsilon}\sum_{N_1}\frac{1}{(q^{1-4\varepsilon}v)^{\deg N_1}}\sum_{N_2}\frac{1}{(q^{1-8\varepsilon}\bar{v})^{\deg N_2}}.
\end{align}
The right-hand side converges in
\begin{align}
    S_3=\{(u,v)||v|>1,|uv^4|<q^{-3}\},
\end{align}
except the possible pole at $u|v^4|q^{2}=q^{-2}$.
Notably, when we specialize to $v=q^{-\frac{1}{2}}$, this pole corresponds to $u=q^{-2}$ and the factor $\frac{1-v}{v^2q-v}=1$. This recovers exadtly the residue previously computed at $u=q^{-2}$.

\subsection{Complete the proof}
We now conclude the proof by synthesizing the convergence information gathered from all regions. The convex hull of $S_1$, $S_{2,1}$, $S_{2,2}$ and $S_3$ is 
\begin{align}
    S_4=\{(u,v)||u|<q^{-1}, |uv^4|^{-3}\}.
\end{align}
By Theorem~\ref{Bochner} and Theorem~\ref{Extending inequalities}, $(u-q^{-2})A_2(u,v)$ has a holomorphic continuation to $S_4$.

We begin by considering the case where $\Re(s)$ and $\Re(w)$ are sufficiently large, corresponding to $|u|$ and $|v|$ being small. The simple pole $u=q^{-2}$ is contained within $S_4$. The shape of $S_4$ allows us to shift the contour from $|u|=r$ to $|u|=q^{-(1+\varepsilon)}$, while setting $v=q^{-\frac{1}{2}}$. The residue at $u=q^{-2}$ gives the main term of \eqref{main}. For the integration along the contour $|u|=q^{-(1+\varepsilon)}$, we estimate the integral trivially, which leads to an additional error term of size $q^{(\frac{1}{2}+\varepsilon)g}$.

%\section{analytical behavior of $A_1(u,v)$}

  \section*{Acknowledgement}
The authors would like to express their sincere gratitude to Professor Peng Gao for proposing the subject of this study and for providing invaluable guidance and insightful recommendations throughout the research process.
%\newpage

\bibliographystyle{plain}
\bibliography{ref}

\end{document}